# Telegrapher’s Generative Model via Kac Flows

### v2

@@ 7.2 Image Generation
Now let $R_{t}\coloneqq f(t)X_{0}+B_{g(t)}$ with a standard Brownian motion $B_{t}$, and choose
$f(t)=1-t,\qquad g(t)=t^{2}.$
@@
[…]
Notice that (98) matches exactly the forward process used in recent flow matching works see [17, 41]. We experimented with various schedules and found that the above flow matching pair
@@
yielded good generation results. We compare our approach to those using the process $(R_{t})_{t}$, i.e. flow matching based on the independent coupling.
@@
We test our approach on generating CIFAR-10 images. We train the velocity field using conditional flow matching with the vector field (7.2) and compare a variety of choices for $(a,c)$. We apply a UNET architecture to parameterize the velocity field and use an ODE solver to simulate the corresponding flow ODE. We used the same setup employed in [39] to train our models. Our code is available online [^1] https://github.com/JChemseddine/telegraphers . For further details we refer to the implementation details in Appendix A.2.
In Table 1 we report the FID results for our methods as well as the baselines. Note that we implemented the baseline ourselves and show the resulting FID as "ours", and we show the reported result in [39] as "reported". In Figures 6, 7 we plot a selection of the generated images. Recall again that $a$ corresponds to the damping coefficient and $c$ to the propagation speed. We found that a combination of sufficiently high damping and lower speeds led to the best generation results. With this experiment we aim to demonstrate the scalability of our method, an extensive study of the effect of the parameters $(a,c)$ is not in the scope of the paper and left to future work.
Table 1: FID scores (100-step Euler sampling) for various $(a,c)$-Kac settings, plus the flow matching baseline.
@@
$a=900,\,c=10$ 18.08 $a=25,\,c=1$ 8.71
$a=900,\,c=20$ 43.39 $a=25,\,c=2$ 6.80
$a=900,\,c=30$ 67.97 $a=25,\,c=3$ 7.64
$a=100,\,c=1$ 12.12 $a=25,\,c=4$ 10.75
$a=100,\,c=3$ 10.78 $a=25,\,c=5$ 16.79
$a=100,\,c=5$ 17.31 FM (reported) 4.64
$a=100,\,c=7$ 26.97 FM (ours) 7.59
$a=100,\,c=10$ 41.34
@@
FM
Figure 6: Generation results of the best performing mean-reverting Kac models and the baseline FM model.
$(a,c)=(100,1)$
$(a,c)=(100,3)$
$(a,c)=(900,10)$
Figure 7: Further generation results of the mean-reverting Kac models.
@@ A.1 Simulating the Kac Process
Recall that the Kac process $K_{t}$ starting in 00 is given by
@@ A.2 Implementation Details
i.e. we draw $X_{0}$ directly from our GMM and add the exact Kac noise $K_{t}$, thereby avoiding any sampling approximation in this toy example. Analogously, in the diffusion case, we use a standard Brownian motion $B_{t}$ instead of $K_{t}$. All models were trained for $500{,}000$ iterations with a learning rate of $5\times 10^{-4}$ and a batch size of $256$. For validation we compute the negative log-likelihood (NLL) of $N$ model-generated samples under the true GMM density:
@@
where $x^{(n)}$ are drawn from the learned model. We set $N=5000$. For the velocity field we used a fully-connected MLP with three hidden layers of width $w=256$ and ReLU activations. Specifically for the diffusion model we required time truncation, namely we train the flow on the interval $[\varepsilon,\,T]$ with $\varepsilon=10^{-15}$. We used the torchdiffeq [6] package to simulate the ODEs with "dopri5" adaptive step size solver.
@@
We trained all models for $400$k iterations, a batch size of $128$ and a learning rate of $2\times 10^{-4}$. We used the torchdiffeq [6] package to simulate the ODEs with an explicit Euler scheme. We use the time dependent U-Net architecture from [26]. We use the same network parameters as in [39], also employing an EMA with decay rate $0.9999$. Both training and evaluation was done using fixed random seeds. The FID is computed on $50$k samples using the torch fidelity package [29]. The code is available at https://github.com/JChemseddine/telegraphers.

### v3

@@ 7.2 Image Generation
Now let $R_{t}\coloneqq f(t)X_{0}+B_{g(t)}$ with a standard Brownian motion $B_{t}$. If we choose
$f(t)=1-t,\qquad g(t)=t^{2},$
@@
$R_{t}=(1-t)\,X_{0}+B_{t^{2}}\overset{d}{=}(1-t)\,X_{0}+t\,B_{1},\quad t\in[0,1].$ (98)
Notice that (98) matches exactly the forward process used in recent flow matching works, see [19, 43]. We experimented with various schedules and found that the above flow matching pair
@@
also yielded good generation results for our Kac model. In the following, we compare our approach via $(M_{t})_{t}$ with that using the process $(R_{t})_{t}$, in particular with flow matching based on the independent coupling.
@@
We test our approach on generating CIFAR-10 images. We train the velocity field using conditional flow matching with the vector field (7.2) and compare a variety of choices for $(a,c)$ and the schedules $g(t)=t$ and $g(t)=t^{2}$. We apply a UNET architecture to parameterize the velocity field and use an ODE solver to simulate the corresponding flow ODE. We aligned our hyperparameters with those employed in [41] to train our models. Our code is available online [^1] https://github.com/JChemseddine/telegraphers. For further details we refer to the implementation details in Appendix A.2.
In Table 1 we report the FID results of our Kac methods via $M_{t}$ as well as the diffusion baselines via $R_{t}$. We implemented the baseline models ourselves and denote them by "Diff" for $g(t)=t$ and "FM" for $g(t)=t^{2}$. In Figures 6, 7 we plot a selection of the generated images. Recall that the parameter $a$ corresponds to the damping coefficient and $c$ to the propagation speed. The schedule $g(t)=t$ generally led to worse performance, especially for the baseline and our methods with diffusion-like parameters. However certain combinations of high damping and low velocity pairs $(a,c)$ such as $(25,2)$ were able to perform well. In contrast the schedule $g(t)=t^{2}$ improved the baseline substantially, and our methods across the board. With this experiment we aim to demonstrate the scalability of our method, an extensive study of the effect of the parameters $(a,c)$ is not in the scope of the paper and left to future work. For further visual examples see Appendix A.3.
Schedule: $g(t)=t^{2}$
@@
$a=900,c=10$ 6.84 $a=100,c=10$ 8.15
$a=900,c=20$ 7.24 $a=25,c=1$ 8.27
$a=900,c=30$ 7.58 $a=25,c=2$ 7.47
$a=100,c=1$ 11.27 $a=25,c=3$ 7.32
$a=100,c=3$ 7.16 $a=25,c=4$ 9.73
$a=100,c=5$ 7.20 $a=25,c=5$ 15.26
$a=100,c=7$ 7.52 FM (our impl.) 7.59
Schedule: $g(t)=t$
Method FID Method FID
$a=900,c=10$ 18.07 $a=100,c=10$ 41.34
$a=900,c=20$ 43.39 $a=25,c=1$ 8.71
$a=900,c=30$ 67.98 $a=25,c=2$ 6.80
$a=100,c=1$ 12.12 $a=25,c=3$ 7.64
$a=100,c=3$ 10.78 $a=25,c=4$ 10.75
$a=100,c=5$ 17.31 $a=25,c=5$ 16.79
$a=100,c=7$ 26.97 Diff (our impl.) 82.27
Table 1: FID (100-step Euler) for $(a,c)$-Kac under two time schedules.
$(a,c)=(900,10)$
$(a,c)=(100,3)$
FM
Figure 6: Generation results of the best performing mean-reverting Kac models and the diffusion baseline (flow matching) for $g(t)=t^{2}$.
@@
Diff
Figure 7: Generation results of the best performing mean-reverting Kac models and the diffusion baseline for $g(t)=t$.
@@ A.1 Simulating the Kac Process
Recall that the Kac process $K_{t}$ starting in 0 is given by
@@ A.2 Implementation Details
i.e. we draw $X_{0}$ directly from our GMM and add the exact Kac noise $K_{t}$, thereby avoiding any sampling approximation in this toy example. Analogously, in the diffusion case, we use a standard Brownian motion $B_{t}$ instead of $K_{t}$. All models were trained for 500,000500{,}000 iterations with a learning rate of $5\times 10^{-4}$ and a batch size of $256$. For validation we compute the negative log-likelihood (NLL) of $N$ model-generated samples under the true GMM density:
@@
where $x^{(n)}$ are drawn from the learned model. We set $N=5000$. For the velocity field we used a fully-connected MLP with three hidden layers of width $w=256$ and ReLU activations. Specifically for the diffusion model we required time truncation, namely we train the flow on the interval $\varepsilon,\,T]$ with $\varepsilon=10^{-15}$. We used the torchdiffeq [8] package to simulate the ODEs with "dopri5" adaptive step size solver.
@@
We trained all models for 400400k iterations, a batch size of $128$ and a constant learning rate of $2\times 10^{-4}$. We used the torchdiffeq [8] package to simulate the ODEs with an explicit Euler scheme. We use the time dependent U-Net architecture from [28]. We use the same network parameters as in [41], also employing an EMA with decay rate $0.9999$. We do not use any data augmentation for training. Note that in our implementation of the diffusion model with schedule $g(t)=t$, we use the common time truncation of training on the interval $\varepsilon,1]$ with $\varepsilon=10^{-5}$ (see [37]). Both training and evaluation was done using fixed random seeds. The FID is computed on 5050k samples using the torch fidelity package [31]. The code is available at https://github.com/JChemseddine/telegraphers.
A.3 Further numerical examples
Here we showcase the CIFAR-10 images generated by our Kac model for other damping/velocity pairs $(a,c)$. The corresponding FID scores can be found in Table 1.
$(a,c)=(25,3)$
$(a,c)=(100,5)$
$(a,c)=(900,20)$
Figure 10: Further generation results of the mean-reverting Kac models with $g(t)=t^{2}$.
$(a,c)=(100,1)$
$(a,c)=(100,3)$
$(a,c)=(900,10)$
Figure 11: Further generation results of the mean-reverting Kac models with $g(t)=t$.